# Sijia Liu2,3, Mingyi Hong1

### v1

@@ Abstract
Unlearning in large language models (LLMs) is critical for regulatory compliance and for building ethical generative AI systems that avoid producing private, toxic, illegal, or copyrighted content. Despite rapid progress, in this work we show that almost all existing unlearning methods fail to achieve true forgetting in practice. Specifically, while evaluations of these ‘unlearned’ models under deterministic (greedy) decoding often suggest successful knowledge removal using standard benchmarks (as has been done in the literature), we show that sensitive information reliably resurfaces when models are sampled with standard probabilistic decoding. To rigorously capture this vulnerability, we introduce leak@$k$, a new meta-evaluation metric that quantifies the likelihood of forgotten knowledge reappearing when generating $k$ samples from the model under realistic decoding strategies. Using three widely adopted benchmarks, TOFU, MUSE, and WMDP, we conduct the first large-scale, systematic study of unlearning reliability using our newly defined leak@$k$ metric. Our findings demonstrate that knowledge leakage persists across methods and tasks, underscoring that current state-of-the-art unlearning techniques provide only limited forgetting and highlighting the urgent need for more robust approaches to LLM unlearning.
footnotetext: $\ast$ Equal contribution.
@@ 1 Introduction
Large language models (LLMs) have demonstrated an extraordinary ability to generate human-like text touvron2023llama. These models are typically pre-trained and fine-tuned on massive datasets collected from the web. However, such datasets often contain harmful, toxic, private, or copyrighted content. This raises significant privacy and ethical concerns, as LLMs may produce biased kotek2023gender; motoki2023more, toxic, private, or illegal responses nasr2023scalable; wen2023unveiling; karamolegkou2023copyright; sun2024trustllm, and even provide dangerous guidance on developing bioweapons or conducting cyberattacks barrett2023identifying; li2024wmdp. To address these risks, LLM unlearning has emerged as a promising approach: the goal is to remove undesired knowledge and its downstream effects while preserving overall model utility.
Unlearning Algorithms. A growing body of work has proposed different unlearning algorithms, often formulating the task as a trade-off between forgetting targeted information and retaining useful capabilities. Examples include gradient ascent methods maini2024tofu, negative preference optimization (NPO) zhang2024negative, simplified NPO variants (SimNPO) fan2024simplicity, representation misdirection (RMU) li2024wmdp, and bi-level or multi-task optimization approaches reisizadeh2025blur; bu2024unlearning. These methods achieve partial success in mitigating unwanted information while preserving model utility. Most approaches rely on supervised unlearning (SU) with token-level cross-entropy loss (see Appendix C), where the model is trained to assign minimum probability to the reference token at each step. This training strategy enforces behavior aligned with the reference outputs. Conversely, RMU li2024wmdp follows an unsupervised strategy where instead of using labeled reference tokens, it modifies hidden representations to shift the model away from the forget set while aiming to maintain performance on the retain set.
Benchmarks for Unlearning. Alongside these algorithmic advances, several benchmarks have been introduced to evaluate unlearning performance, such as TOFU maini2024tofu, MUSE shi2024muse, WMDP li2024wmdp, and the multi-task benchmark LUME ramakrishna2025lume. These benchmarks test whether models avoid reproducing sensitive knowledge while continuing to generate accurate and useful outputs on non-forget tasks (see Appendix A for details).
@@
Challenges. A fundamental limitation in the current unlearning literature lies in the mismatch between evaluation and deployment settings. Nearly all existing benchmarks rely on deterministic decoding, most commonly greedy decoding, where the model always selects the single most probable token at each step. While this setup is convenient for standardized evaluation, it poorly reflects deployed systems, where probabilistic decoding strategies such as temperature sampling or top-$p$ sampling are widely adopted, especially in domains such as conversational agent holtzman2019curious; chung2023increasing and code generation chen2021evaluating; arora2024optimizing.
This reliance on greedy decoding creates a serious blind spot. Greedy outputs are low-diversity and repetitive, and often underestimate the model’s latent knowledge. In contrast, probabilistic decoding exhibits creativity  nguyen2024turning and generates human-preferred text holtzman2019curious, but it also raises the risk that suppressed knowledge resurfaces if undesired knowledge is not truly forgotten  krishnan2025not; scholten2024probabilistic. In the unlearning tasks, even a single leaked generation can be catastrophic, for example exposing copyrighted passages or private information.
@@
Table 1: Examples of generated text from different unlearning methods on the MUSE-News dataset, comparing outputs from the deterministic decoder with the worst-case response among $64$ generations from the probabilistic decoder with $T=0.2$ and $p=1.0$. Failed unlearning is indicated by undesired answers in red, successful unlearning by green.
For example, as Table 1 illustrates, models that appear to have forgotten sensitive passages under greedy decoding readily regenerate them verbatim once sampled multiple times under the probabilistic decoding. In particular, it presents an example from MUSE-News benchmark using the LLaMA2-7B model. Several unlearning methods, including GradDiff liu2022continual, NPO zhang2024negative, and BLUR-NPO reisizadeh2025blur are evaluated under two decoding strategies: (1) a greedy decoder, as typically used in current benchmarks, and (2) a probabilistic decoder, where we report the most privacy-leaking response among 64 generations, setting T=0.2T\!=\!0.2 and p=1.0p\!=\!1.0. Hence, existing benchmark results on unlearned models may misleadingly determine that forgetting has succeeded, when in fact this “success” is brittle and fails to generalize to realistic usage.
Most unlearning evaluations adopt greedy decoding due to its deterministic and reproducible results, while only a few consider probabilistic decoding. Notably, scholten2024probabilistic and yuan2024closer first explore probabilistic evaluation. However, their approaches remain limited: (1) they measure only statistical uncertainty in token distributions, without assessing whether the generated outputs convey the semantic content of the forgotten knowledge; (2) they rely on single-generation evaluation, which is problematic because one sample provides only a narrow view of the model’s output space and can easily miss residual traces of forgotten knowledge that persist under probabilistic decoding. Such discrepancies reveal that current unlearning methods often provide only an illusion of forgetting, undermining claims of privacy, copyright compliance, and safety.
@@ 1.1 Our Contributions
We show in this work that current approaches to LLM unlearning provide only an illusion of forgetting. While prior evaluations suggest that harmful, private, or copyrighted content has been erased, we show that such content readily resurfaces once models are queried under realistic conditions. Specifically, we demonstrate that when sampling with non-zero $T$ or $p$, where $T$ is the decoding temperature, and $p$ is the top-$p$ value, unlearned models continue to leak sensitive knowledge. More concretely, our contributions are listed below:
@@
(1) We introduce leak@$k$, a meta-metric that quantifies the likelihood of sensitive content reappearing after LLMs generate $k$ responses for the same question. Unlike prior evaluation protocols, which rely exclusively on deterministic decoding and therefore underestimate residual memorization, leak@$k$ directly measures the probability that at least one sampled generation reveals targeted information, as determined by core evaluation metrics such as ROUGE-L lin2004rouge, Cosine Similarity reimers-2019-sentence-bert, Entailment Score yuan2024closer, or Accuracy. We provide two unbiased estimators, $\widehat{\texttt{leak@$k$}}$ and $\widehat{\texttt{L}}_{\text{worst-}k}$, of leak@$k$, where the former has lower variance while the latter is relatively easier to implement.
@@
Figure 1: $\widehat{\texttt{leak@$k$}}$ measure using ROUGE-L score ($\widehat{\text{leak@$k$}}$–RS) for various unlearned models on MUSE-News dataset using LLaMA2-7B model at T=0.2T\!=\!0.2 and p=1.0p\!=\!1.0. When $k$ is small, the unlearned models show limited leakage in providing information from the forget set. However, as $k$ increases, all models reveal increasingly sensitive information about the forget set questions.
@@
(2) We conduct the first large-scale systematic study of unlearning reliability under probabilistic decoding. Our experiments cover three widely used benchmarks, TOFU, MUSE, and WMDP maini2024tofu; shi2024muse; li2024wmdp, and evaluate leading unlearning methods across multiple settings of temperature $T$ and top-$p$ sampling. Across almost all settings, our results are strikingly consistent: our meta-metric leak@$k$ rises sharply with $k$, i.e., as more generations are sampled under probabilistic decoding, the probability of producing at least one leaking output rises rapidly (see, e.g., Fig. 1).
@@
In summary, our findings reveal a critical gap between existing evaluation protocols and practical deployment: what appears to be successful forgetting under deterministic decoding often proves weak and unreliable in practice. This highlights an urgent need for new unlearning methods that remain robust under realistic probabilistic decoding, as well as the development of stronger benchmarks that can reliably capture such challenges.
@@
In this section, we introduce our meta-metric leak@$k$, which quantifies the extent of information leakage by evaluating the expected leakage of the most revealing response among $k$ generations for a given prompt. To instantiate this meta-metric, we leverage several core evaluation metrics that are widely adopted in the unlearning literature. Specifically, we discuss ROUGE-L, Cosine Similarity, Entailment Score, Accuracy, and LLM-as-Judge, and further elaborate on how these metrics are integrated to formulate our meta-metric leak@$k$.
ROUGE-L Score (RS) measures the word-level overlap between the model’s generated response $f(q;\bm{\theta})$ to a question $q$ and the corresponding gold answer lin2004rouge.
Cosine Similarity (CS) measures semantic similarity between the generated response $f(q;\bm{\theta})$ and the gold answer $a$ by comparing their contextual embeddings. We compute embeddings using a pretrained sentence transformer model (e.g., Sentence-BERT reimers-2019-sentence-bert) and report the cosine of the angle between the two embedding vectors. The score ranges from 1-1 to $1$, with higher values indicating stronger semantic alignment between $f(q;\bm{\theta})$ and $a$.
Entailment Score (ES) quantifies factual correctness by checking whether a generated answer $f(q;\bm{\theta})$ entails the ground truth $a$, using a pretrained NLI model sileo2023tasksource: $f(q;\bm{\theta})$ is considered to entail $a$ if a human reading the generated answer would typically infer that the gold answer $a$ is most likely true yuan2024closer. The score is binary ($1$ if entailed, 0 otherwise).
@@
LLM-as-Judge (LJ) leverages large language model as an automatic evaluator gu2024survey to determine whether a generated response $f(q;\bm{\theta})$ reveals sensitive information related to the gold answer $a$. The evaluation prompt includes the question $q$, the gold answer $a$, and the generated response $f(q;\bm{\theta})$, and the judge model outputs a score reflecting the correctness of the response with respect to the gold answer. Details of the prompt design and illustrative examples are provided in Appendix B.1. Recently, this framework has also been adopted for unlearning evaluation as a reliable judge wang2025reasoning. For the TOFU dataset, we design the prompt such that the LLM judge provides a binary score (0 or 1), reflecting whether the model successfully forgot the target knowledge. In contrast, for the more complex WMDP benchmark involving reasoning and domain-specific hazardous knowledge, we design the prompt for the LLM judge to provide a graded score from 1 to 3, capturing different levels of knowledge leakage.
Current evaluation metrics provide useful insights into leakage after unlearning but suffer from serious limitations: Most rely on greedy decoding maini2024tofu; shi2024muse; li2024wmdp, which ignores the probabilistic nature of LLMs. Recent work has explored entropy-based probabilistic evaluation scholten2024probabilistic; yuan2024closer, but these approaches focus only on statistical uncertainty and do not capture task-level semantics.
We introduce leak@$k$, a semantic and distributional meta-metric that quantifies the expected leakage of the most leaking response among $k$ generations. As a meta-metric, leak@$k$ can be instantiated with different core metrics (e.g., RS), making it flexible and broadly applicable. To introduce our proposed metric, let us assume that the model generates multiple responses for each question $q$ using a probabilistic decoder. For each response $f(q;\bm{\theta})$, we compute the correctness score $S(q)\!:=\!\sf{CoreM}(a,f(q;\bm{\theta}))\!\in\![0,1]$ where $a$ is the ground-truth answer and $\sf{CoreM}(\cdot,\cdot)$ denotes the used core evaluation metric (e.g. RS). Intuitively, $S(q)$ measures how well the generated response matches the reference, with higher values indicating stronger alignment, which on the forget set corresponds to greater information leakage. The metric leak@$k$, is defined as the expected maximum score among $k$ independent draws, given as
$\texttt{leak@$k$}:=\mathbb{E}\!\left[\max_{1\leq j\leq k}S_{j}\right],$
where $S_{1},\dots,S_{k}$ are i.i.d. correctness scores. Using $\mathbb{E}[X]=\int_{0}^{1}\mathop{\rm Pr}\nolimits(X\geq\tau)\,d\tau$, we can write
@@
In practice, to obtain a low-variance estimate of leak@$k$, we generate $n\geq k$ samples per question and apply the unbiased estimator described below. For a fixed threshold $\tau$, let
@@
[…]
In summary, the proposed meta-metric design follows two key principles: (1) We measure unlearning under probabilistic decoding, which reflects real deployment where LLM outputs are sampled rather than deterministically chosen. (2) We focus on the most leaking response among $k$ generations, since ensuring no leakage even in this worst case provides a sufficient condition for unlearning success. In practice, leak@$k$ is applied by sampling $n>k$ responses per prompt under probabilistic decoding, evaluating each with a core metric, and then computing $\widehat{\texttt{leak@$k$}}$ from these scores to quantify worst-case leakage.

### v2

@@ Abstract
Unlearning in large language models (LLMs) is critical for regulatory compliance and for building ethical generative AI systems that avoid producing private, toxic, illegal, or copyrighted content. Despite rapid progress, in this work we show that almost all existing unlearning methods fail to achieve true forgetting in practice. Specifically, while evaluations of these ‘unlearned’ models under deterministic (greedy) decoding often suggest successful knowledge removal using standard benchmarks (as has been done in the literature), we show that sensitive information reliably resurfaces when models are sampled with standard probabilistic decoding. To rigorously capture this vulnerability, we introduce leak@$k$, a new meta-evaluation metric that quantifies the likelihood of forgotten knowledge reappearing when generating $k$ samples from the model under realistic decoding strategies. Using three widely adopted benchmarks, TOFU, MUSE, and WMDP, we conduct the first large-scale, systematic study of unlearning reliability using our newly defined leak@$k$ metric. Our findings demonstrate that knowledge leakage persists across methods and tasks, underscoring that current state-of-the-art unlearning techniques provide only limited forgetting and highlighting the urgent need for more robust approaches to LLM unlearning. We propose an algorithm, termed Robust Unlearning under LEak@$k$ metric (RULE), which serves as an initial step toward addressing this concern. We demonstrate that RULE provides an unlearned model for TOFU benchmark with no information leakage for a large number of generation samples.
11footnotetext: Equal contribution.
@@ 1 Introduction
Large language models (LLMs) have demonstrated an extraordinary ability to generate human-like text Touvron et al. (2023). These models are typically pre-trained and fine-tuned on massive datasets collected from the web. However, such datasets often contain harmful, toxic, private, or copyrighted content. This raises significant privacy and ethical concerns, as LLMs may produce biased Kotek et al. (2023); Motoki et al. (2023), toxic, private, or illegal responses Nasr et al. (2023); Wen et al. (2023); Karamolegkou et al. (2023); Sun et al. (2024), and even provide dangerous guidance on developing bioweapons or conducting cyberattacks Barrett et al. (2023); Li et al. (2024). To address these risks, LLM unlearning has emerged as a promising approach: the goal is to remove undesired knowledge and its downstream effects while preserving overall model utility.
Unlearning Algorithms. A growing body of work has proposed different unlearning algorithms, often formulating the task as a trade-off between forgetting targeted information and retaining useful capabilities. Examples include gradient ascent methods Maini et al. (2024), negative preference optimization (NPO) Zhang et al. (2024a), simplified NPO variants (SimNPO) Fan et al. (2024), representation misdirection (RMU) Li et al. (2024), and bi-level or multi-task optimization approaches Reisizadeh et al. (2025); Bu et al. (2024). These methods achieve partial success in mitigating unwanted information while preserving model utility. Most approaches rely on supervised fine-tuning (SFT) with token-level cross-entropy loss (see Appendix C), where the model is trained to assign maximum probability to the reference token at each step. This training strategy enforces behavior aligned with the reference outputs. Conversely, RMU Li et al. (2024) follows an unsupervised strategy where instead of using labeled reference tokens, it modifies hidden representations to shift the model away from the forget set while aiming to maintain performance on the retain set.
Benchmarks for Unlearning. Alongside these algorithmic advances, several benchmarks have been introduced to evaluate unlearning performance, such as TOFU Maini et al. (2024), MUSE Shi et al. (2024), WMDP Li et al. (2024), and the multi-task benchmark LUME Ramakrishna et al. (2025). These benchmarks test whether models avoid reproducing sensitive knowledge while continuing to generate accurate and useful outputs on non-forget tasks (see Appendix A for details).
@@
Challenges. Current benchmarks almost exclusively rely on greedy decoding, creating a significant mismatch between evaluation protocols and real-world deployment. While greedy decoding facilitates convenient evaluation and fair comparison for different methods, it poorly reflects deployed systems, where probabilistic decoding strategies such as temperature sampling or nucleus sampling are widely adopted, especially in domains such as conversational agent Holtzman et al. (2019); Chung et al. (2023) and code generation Chen et al. (2021); Arora et al. (2024).
Moreover, this reliance on greedy decoding creates a serious security blind spot. Greedy outputs are low-diversity and repetitive, and failing to surface model’s latent knowledge. In contrast, probabilistic decoding exhibits creativity  Nguyen et al. (2024) and generates human-preferred text Holtzman et al. (2019), but it also raises the risk that suppressed knowledge resurfaces if undesired knowledge is not truly forgotten  Krishnan et al. (2025); Scholten et al. (2024).
@@
Table 1: Examples of generated text on the MUSE-News dataset across three unlearning methods: GradDiff Liu et al. (2022), NPO Zhang et al. (2024a), and BLUR-NPO Reisizadeh et al. (2025). The left column displays results from greedy decoding, while the right column shows the worst-case response among 64 probabilistic generations ($T=0.2,p=1.0$). Failed unlearning is highlighted in red, and successful unlearning in green.
For example, as Table 1 illustrates, the model that appears to have forgotten sensitive passages under greedy decoding readily regenerate them verbatim once sampled multiple times under the probabilistic decoding. Fig. 1 provides one justification for information leakage phenomenon where probabilistic decoding can sample an early sensitive token that greedy decoding consistently suppresses; once this first sensitive token is generated, the model’s conditional distribution over subsequent tokens becomes sharply peaked, causing the remaining sensitive content to be produced with high confidence, more details in Appendix E.2. This property makes probabilistic decoding suitable for reliable evaluation, especially in unlearning tasks where even a single leaked generation can be catastrophic.
Figure 1: Evolution of token prediction probability distributions during decoding. Sampling the initial sensitive token (“1”) significantly increases model confidence for subsequent sensitive tokens (e.g., “3”, “years”).
However, most unlearning evaluations still rely on greedy decoding for its determinism and reproducibility, with limited consideration of probabilistic decoding. Notably, Scholten et al. (2024) takes an initial step toward probabilistic evaluation, but their evaluation has several limitations. First, the analysis remains focused on single-generation behavior: although multiple samples are drawn, leakage is measured per generation and averaged, rather than modeling how leakage escalates across repeated sampling attempts, which can be exploited by adversarial users. Second, their evaluation is limited to token-level distributional metrics and does not assess whether sampled outputs semantically recover the forgotten knowledge. To mitigate leakage, they propose an entropy-regularized objective (NPO+ENT) that reduces output diversity by making the model more deterministic; however, effective unlearning should preserve output diversity while ensuring that sensitive tokens are never sampled. A more detailed discussion is provided in the Related Work in Appendix  A.
@@ 1.1 Our Contributions
We show in this work that current approaches to LLM unlearning provide only an illusion of forgetting. While prior evaluations suggest that harmful, private, or copyrighted content has been erased, we show that such content readily resurfaces once models are queried under realistic conditions. Specifically, we demonstrate that when sampling with non-zero $T$ or $p$, where $T$ is the decoding temperature, and $p$ is the top-$p$ value, unlearned models continue to leak sensitive knowledge. To address this issue, we introduce a robust unlearning method under a multi-sampling setting as an initial step toward robustness under realistic conditions. More concretely, our contributions are listed below:
@@
(1) We introduce leak@$k$, a meta-metric that quantifies the likelihood of sensitive content reappearing after LLMs generate $k$ responses for the same question. Unlike prior evaluation protocols, which rely exclusively on deterministic decoding and therefore underestimate residual memorization, leak@$k$ directly measures the probability that at least one sampled generation reveals targeted information, as determined by core evaluation metrics such as ROUGE-L Lin (2004b), Cosine Similarity Reimers and Gurevych (2019), Entailment Score Yuan et al. (2024), or Accuracy. We provide two unbiased estimators of leak@$k$, namely $\widehat{\texttt{leak@$k$}}$ and $\widehat{\texttt{L}}_{\text{worst-}k}$, where the former has lower variance while the latter is relatively easier to implement.
@@
Figure 2: $\widehat{\texttt{leak@$k$}}$ measure using entailment score ($\widehat{\text{leak@$k$}}$–ES) for various unlearned models on TOFU dataset using LLaMA-3.2-1B model at T=1.0T\!=\!1.0 and p=1.0p\!=\!1.0. As $k$ increases, all of the previous methods reveal increasingly sensitive information about the forget set questions. Our algorithm RULE–GradDiff achieves almost $ES=0$ across all values of $k$.
@@
(2) We conduct the first large-scale systematic study of unlearning reliability under probabilistic decoding. Our experiments cover three widely used benchmarks, TOFU, MUSE-News, and WMDP Maini et al. (2024); Shi et al. (2024); Li et al. (2024), and evaluate leading unlearning methods across multiple settings of temperature $T$ and top-$p$ sampling. Across almost all settings, our results are strikingly consistent: our meta-metric leak@$k$ rises sharply with $k$, meaning that as more generations are sampled under probabilistic decoding, the probability of producing at least one leaking output rises rapidly. see, e.g., Fig. 2. Across almost all benchmarks and unlearning methods, the performance rankings are preserved as $T$, $p$, and $k$ change and are consistent with the ordering obtained under greedy decoding.
(3) We propose a new algorithm named Robust Unlearning under LEak@$k$ metric (RULE) for QA based task. At a high level, in this algorithm, we adjust the forget set dynamically at each unlearning epoch. During each training epoch, we prompt the (current) model multiple times using a fixed query from the original (forget) set and the probabilistic decoding. Then, we enhance the forget set with the most leaking responses that are identified by exploiting an evaluation metric. Finally, we run the optimization step using the enhanced forget set. RULE demonstrates an outstanding performance on the TOFU benchmark achieving almost $ES=0$ (NO information leakage) even we prompt the unlearned model $128$ times.
@@
In summary, our findings reveal a critical gap between existing evaluation protocols and practical deployment: what appears to be successful forgetting under deterministic decoding often proves weak and unreliable in practice. Moreover, RULE serves as the initial step toward a robust unlearning approach under leak@$k$ metric.
@@ 2 Leak@$k$: A Meta-Metric for Reliable Unlearning Evaluation
In this section, we introduce our meta-metric leak@$k$, which quantifies the extent of information leakage by evaluating the expected leakage of the most revealing response among $k$ generations for a given prompt. Our meta-metric can be instantiated with a specific core evaluation metric. Specifically, we discuss ROUGE-L Score, Cosine Similarity, Entailment Score, Accuracy, and LLM-as-Judge that are widely adopted in the unlearning literature. We further elaborate on how these base metrics work, and how to integrate them to formulate our meta-metric leak@$k$.
ROUGE-L Score (RS) measures the word-level overlap between the model’s generated response $f(q;\bm{\theta})$ to a question $q$ and the corresponding gold answer Lin (2004b).
Cosine Similarity (CS) measures semantic similarity between the generated response $f(q;\bm{\theta})$ and the gold answer $a$ by comparing their contextual embeddings. We compute embeddings using a pretrained sentence transformer (e.g., Sentence-BERT Reimers and Gurevych (2019)) and report the cosine of the angle between the two embedding vectors, which takes value in (1,1)(-1,1), with higher values indicating stronger semantic alignment between $f(q;\bm{\theta})$ and $a$.
Entailment Score (ES) quantifies factual correctness by checking whether a generated answer $f(q;\bm{\theta})$ entails the ground truth $a$, using a pretrained NLI model Sileo (2023): $f(q;\bm{\theta})$ is considered to entail $a$ if a human reading the generated answer would typically infer that the gold answer $a$ is most likely true Yuan et al. (2024). The score is binary ($1$ if entailed, 0 otherwise).
@@
LLM-as-Judge (LJ) leverages large language model as an automatic evaluator Gu et al. (2024) to determine whether a generated response $f(q;\bm{\theta})$ reveals sensitive information related to the gold answer $a$. The evaluation prompt includes the question $q$, the gold answer $a$, and the generated response $f(q;\bm{\theta})$, and the judge model outputs a score reflecting the correctness of the response with respect to the gold answer. Details of the prompt design and illustrative examples are provided in Appendix B.1. Recently, this framework has also been adopted for unlearning evaluation as a reliable judge Wang et al. (2025). For the TOFU dataset, we design the prompt such that the LLM judge provides a binary score (0 or 1), reflecting whether the model successfully forgot the target knowledge. In contrast, for the more complex WMDP benchmark involving reasoning and domain-specific hazardous knowledge, we design the prompt for the LLM judge to provide a graded score from 1 to 3, capturing different levels of knowledge leakage.
Current evaluation metrics provide useful insights into leakage after unlearning but suffer from serious limitations: Most rely on greedy decoding Maini et al. (2024); Shi et al. (2024); Li et al. (2024), which ignores the probabilistic nature of LLMs. Recent work has explored entropy-based probabilistic evaluation Scholten et al. (2024); Yuan et al. (2024), but these approaches focus only on statistical uncertainty and do not capture task-level semantics. Detailed discussion in Appendix A.
We introduce leak@$k$, a distributional meta-metric that quantifies the expected leakage of the most leaking response among $k$ generations. As a meta-metric, leak@$k$ can be instantiated with different core metrics (e.g., RS), making it flexible and broadly applicable. To introduce our proposed metric, let us assume that the model generates multiple responses for each question $q$ using a probabilistic decoder. For each response $f(q;\bm{\theta})$, we compute the correctness score $S(q)\!:=\!\sf{CoreM}(a,f(q;\bm{\theta}))\!\in\![0,1]$ where $a$ is the ground-truth answer and $\sf{CoreM}(\cdot,\cdot)$ denotes the used core evaluation metric (e.g. RS). Intuitively, $S(q)$ measures how well the generated response matches the reference, with higher values indicating stronger alignment, which on the forget set corresponds to greater information leakage. The metric leak@$k$, is defined as the expected maximum score among $k$ independent draws, given as $\texttt{leak@$k$}:=\mathbb{E}\!\left[\max_{1\leq j\leq k}S_{j}\right]$ where $S_{1},\dots,S_{k}$ are i.i.d. correctness scores. Using $\mathbb{E}[X]=\int_{0}^{1}\mathop{\rm Pr}\nolimits(X\geq\tau)\,d\tau$, we can write
@@
In practice, to get a low-variance estimate of leak@$k$, we generate $n\geq k$ samples per question and apply the unbiased estimator described below. For a fixed threshold $\tau$, let
@@
In summary, the proposed meta-metric design follows two key principles: (1) We measure unlearning under probabilistic decoding, which reflects real deployment where LLM outputs are sampled rather than deterministically chosen; (2) We focus on the most leaking response among $k$ generations, since ensuring no leakage even in this worst case provides a sufficient condition for unlearning success. In practice, leak@$k$ is applied by sampling $n>k$ responses per prompt under probabilistic decoding, evaluating each with a core metric, and then computing $\widehat{\texttt{leak@$k$}}$ from these scores to quantify worst-case leakage.